OpenStudy (anonymous):

assume linear depreciation: Let x denote a temperature on the celsius scale, and let y denote the corresponding temperature on the fahrenheit scale. a) find a linear function relating x and use the facts that 32 degrees F corresponds to 0c and 212F corresponds to 100 degrees. write the function in the form y=Ax+B. b) what celsius temperature corresponds to 98.6 F? c) find a number z for which xdegreeF=zdegreeC.

OpenStudy (anonymous):

I have no clue any suggestions

OpenStudy (saifoo.khan):

Lol.

OpenStudy (saifoo.khan):

Mini will give some.

OpenStudy (anonymous):

I thought now correct me if I am wrong (32, 0) and (212, 100) am I on the right track

OpenStudy (mimi_x3):

Saifi, you can try first :D

OpenStudy (saifoo.khan):

Ladies first. @mini

OpenStudy (anonymous):

this lady is clueless

OpenStudy (anonymous):

I am and Im a lady

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

LOl @mini. should i take your name over here as well?

OpenStudy (saifoo.khan):

Ohh i thought u were male @danyel.

OpenStudy (mimi_x3):

You can use the censored version :P Danyel is a guys name >_<

OpenStudy (saifoo.khan):

No, i will use uncensored.. since u are doing the same.

OpenStudy (anonymous):

its ok a lot of people do I just spell it different you see it (Danielle) most of the time.

OpenStudy (mimi_x3):

Sorry, deleted it :D Nice name btw

OpenStudy (saifoo.khan):

Wow, code words.

OpenStudy (anonymous):

you are funny Its ok

OpenStudy (saifoo.khan):

Lol.

OpenStudy (saifoo.khan):

btw @danyel.. u were correct.

OpenStudy (anonymous):

ok I was hoping

OpenStudy (anonymous):

how do I plug that into y=Ax+B

OpenStudy (saifoo.khan):

slope, then y interceot.

OpenStudy (saifoo.khan):

intercept*

OpenStudy (anonymous):

am i doing the slope = (100-0)/(212-32)

OpenStudy (saifoo.khan):

right,

OpenStudy (anonymous):

=100/170

OpenStudy (saifoo.khan):

right. 5/9 slope.

OpenStudy (anonymous):

I was looking for my calc thanks

OpenStudy (saifoo.khan):

np. :)

OpenStudy (anonymous):

now I do y=mx+b

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (saifoo.khan):

i better find a job in colleges as a teacher. LOL.

OpenStudy (anonymous):

no kidding

OpenStudy (anonymous):

so here I get confused what numbers do I stick where?

OpenStudy (saifoo.khan):

u have two cordinates. use any one..

OpenStudy (anonymous):

lets use (32, 0) where do i plug themin at?

OpenStudy (anonymous):

them in

OpenStudy (saifoo.khan):

hmm.. carry on..

OpenStudy (anonymous):

y=1(x/32)+b

OpenStudy (saifoo.khan):

u forgot a 0/.

OpenStudy (anonymous):

ok so it will say (0x/32)+B

OpenStudy (saifoo.khan):

No, lol.

OpenStudy (saifoo.khan):

( x,y) (32,0) y=5x/9+b 0 = 5(32)/9 +b

OpenStudy (anonymous):

0=5(32)/9+B

OpenStudy (saifoo.khan):

yo.

OpenStudy (anonymous):

then i go 160/9+B?

OpenStudy (saifoo.khan):

dont forget zero.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

0=18+b

OpenStudy (saifoo.khan):

how u did that? :o

OpenStudy (anonymous):

I took 160/9

OpenStudy (anonymous):

I don't do that

OpenStudy (saifoo.khan):

no, how u got 18?

OpenStudy (anonymous):

It was 17.777 so i rounded up

OpenStudy (saifoo.khan):

No. i prefer u take that in fraction instead.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

160/9 then

OpenStudy (saifoo.khan):

b = -160/9

OpenStudy (anonymous):

right that is what i got

OpenStudy (saifoo.khan):

hmm. right.

OpenStudy (anonymous):

so that is the slope right

OpenStudy (saifoo.khan):

No. the slope is 5/9. this is the y intercept.

OpenStudy (anonymous):

thats right sorry I am getting confused with so many steps

OpenStudy (saifoo.khan):

i can understand...

OpenStudy (saifoo.khan):

dont worry. ^_^

OpenStudy (anonymous):

This is all the answer to A) correct?

OpenStudy (saifoo.khan):

wait. the answer is \[y = \frac59x - \frac{160}{9}\]

OpenStudy (anonymous):

/I knew what you ment

OpenStudy (saifoo.khan):

i wrote that incorrect,

OpenStudy (anonymous):

so now I have to find out what temperature corresponds to 98.6 degrees F.

OpenStudy (saifoo.khan):

insert the values to get the answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the 98.6

OpenStudy (saifoo.khan):

are u sure? im getting something different.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I can't think any more I need to go to sleep Thanks for helping me tonite

OpenStudy (anonymous):

I can't seem to process what I need to do sorry.

OpenStudy (saifoo.khan):

lol, ok finish this part first then u may go to bed.

OpenStudy (saifoo.khan):

98.6 = (5 x)/9-160/9 solve for x.